OpenStudy (anonymous):

Solve 2|5x − 2| + 4 = 18 for x. Show your work to receive full credit.

OpenStudy (anonymous):

I'm fine with partial credit, so I'll have you do the work for me: Think about your order of operations here, first. Remember, when solving equations, you use the reverse order of operations, so you'll subtract 4 from both sides first. That leaves you with: 2|5x - 2| = 14 You can now divide both sides by two: |5x - 2| = 7 Now, to deal with the absolute value: the inside of the absolute value must be either negative or positive 7. So we now have two equations: 5x - 2 = 7 5x - 2 = -7 Now, simply solve for "x" in both equations! Remember your order of operations, and show full work to receive credit. ;)